# Supplemental Material Appendix A Additional Related Works A.1 Image Denoising and Restoration

### v3

@@ A.1 Image Denoising and Restoration
Image denoising and restoration belong to a set of problems called image inverse problems [Ongie2020-zi]. There are several different levels of difficulty for image denoising and restoration problems. The easiest level is when the corruption process is known as a linear process (such as blurring, inpainting, uniform Gaussian noise, etc) and can be programmatically replicated onto clean images, and the goal is to create an inverse process to the corruption. There have been extensive studies of restoring known corruptions, including supervised approaches [Ongie2020-zi] and unsupervised approaches [Kawar2022-vc, Kadkhodaie2021-wd, wang2022zero]. The second level of difficulty is when the corruption process is not reproducible, but there are existing pairs of clean and corrupted images allowing for supervised training, and the typical solution is training models to generate the clean image given the corrupted images in a supervised way, such as UNet [Ronneberger2015-um, Manifold2019-ft] or Conditional Diffusion [Choi2021-mx, Kawar2022-vc, Saharia2022-co]. The most difficult setting is when we don’t have paired data and also cannot reproduce the corruption programmatically; instead, we only have access to unpaired clean images and corrupted images. Existing approaches for this challenging setting include cycle-consistency-based approaches such as CycleGAN [Zhu2017-wf] and UnitDDPM [Sasaki2021-ks], VAE-based approaches [Zheng2023-dz], and flow-based approaches [Du2020-om]. In addition to the above settings, there are also denoising methods that do not require a training dataset, such as MedianBlur or Deep Image Prior [Ulyanov2016-kz]. General Diffusion Prior [fei2023generative] is a recent unsupervised diffusion-based method that can perform unpaired non-linear blind image restoration, but it is extremely computationally costly as it requires full generative diffusion with multiple gradient steps within each reverse diffusion step.
Restoring noisy SRH images generally falls within the third difficulty level (i.e. no paired data and not a linear inverse problem), as the distribution of noise is unpredictable. Although it is possible to apply supervised methods from the second level by training to remove Gaussian and Laser noises (such as our UNet and Conditional Diffusion baselines), they are generally outperformed by unpaired methods as shown in our experiments. Several recent approaches (including [chung2022come], [Chung2023-qt], and [he2023tdiffde]) also used the last part of the reverse diffusion process for unpaired image restoration. We discussed these closely related works in more detail in Section LABEL:sec:relwork.
@@ A.2 Deep Learning applications on SRH
Deep learning techniques have been playing an important role ever since SRH was used for intra-operative neurosurgery, as the very first paper that introduces this optical diagnostic technique [Orringer2017-nn] trained a multi-layer perceptron to automatically predicts brain tumor subtypes from quantified SRH image attributes. Later, Hollon et al. [Hollon2020-ez] proposed a CNN-based model that can directly interpret raw SRH images (instead of quantified attributes) and make diagnosis at over 94% accuracy. The first public dataset of clinical SRH images, OpenSRH [Jiang2022-cj], was released in 2022, and it contains SRH images from 300+ brain tumor patients and 1300+ unique whole slide optical images. DeepGlioma [Hollon2023-uf] has allowed deep-learning-based models to perform molecular classification of diffuse glioma types via a multimodal model that takes in both SRH images and diffuse glioma genomic data. HiDisc [Jiang2023-rq] and S3L [hou2024selfsupervised] are self-supervised visual representation learning methods that were developed to classify whole-slide SRH images.
All the above works focused on the classification of SRH images, and there have been few previous works that attempted to tackle the unavoidable noise problem of SRH images through image restoration. To the best of our knowledge, the only previous work on this topic was [Manifold2019-ft], in which the authors collected a paired dataset using low and high laser energy, and trained a U-Net to perform the denoising. In comparison, our method works for all types of noisy SRH images (not just those resulting from low laser power), and we have shown that our method outperforms their U-Net approach.
@@ Appendix C Methodology Additional Details
C.1 Step Calibrator Training and Calibration Distribution
We show the full algorithm of training the step calibrator with manually added Gaussian noise and augmentations in Algorithm C.1, as well as one example of the augmentation in Figure 2.
[…]
@@ C.1 Step Calibrator Training and Calibration Distribution
[width=0.6]figs/aug.png
[width=0.6]figs/histogram.png
C.2 Diffusion Model Training
We show the full training algorithm in Algorithm C.2 below. It is almost identical to the training algorithm from [Ho2020-tm], except that we only sample up to .
[…]
{algorithmic}[1]
[…]
@@ D.1 Baseline and Ablation details
CycleGAN [Zhu2017-wf]: We trained a CycleGAN model for image-to-image translation between unpaired low-quality and high-quality images. We used the 4.5K hand-picked high-quality and 3.8K hand-picked low-quality images for training the CyleGAN. Both the high-quality and low-quality image sets contained at least 10 different tumor types. The CycleGAN model performed better when trained on the manually selected images compared to training on the full 840K filtered high-quality images.
Synthetic noise[Lehtinen2018-zd, Manifold2019-ft]: A U-Net model is trained by adding Gaussian noise to high-quality images to generate synthetic low-quality images. We train only with the 4.5K manually selected high-quality images because we found that the model performs better compared to training on the full 840K filtered high-quality images.
Conditional Diffusion [Kawar2022-vc, Saharia2022-co]: We train a conditional generative diffusion model, where we sample from a random Gaussian prior, and generate the restored image conditioned on the low-quality image. The conditioning method is concatenation (i.e. the input image and the conditioning image is concatenated as input to the UNet). We train the model with the 840K HQ training images and fine-tune it with the 4.5K hand-picked images. During training, we add a random amount of Gaussian noise to the high-quality images as conditioning images. We show the full training algorithm in Algorithm.
Conditional Diffusion: Training
{algorithmic}[1]
[…]
Regularized Reverse Diffusion (RRD) [Chung2023-qt] is a recent method developed for restoring noisy medical images. It also uses part of the reverse diffusion process as restoration, but it employs a non-parametric noise estimator [Chen2015-ij] to determine the step numbers, and uses a Fourier-feature-based regularization to reduce hallucination. Since we were unable to find any official code for the RRD paper and we deal with images in a vastly different domain from theirs (RRD was originally designed for denoising single-channel MRI images), we trained our own diffusion model following the same process as [Ho2020-tm] and using the same cosine noise schedule as RSCD.
Deep Image Prior [Ulyanov2016-kz] is an unsupervised image restoration method that works by training a ResNet to reconstruct the noisy image but takes the generation of the model before its convergence as the denoised image. We take the output of step 1800 as our baseline, as images generated from this training step tend to have the best quality.
@@
No Dynamic Recalibration: RSCD without Dynamic Recalibration, i.e. Algorithm LABEL:alg:dr without the blue parts. We take the initial step calibration and directly run that many steps of denoising to the end.
Nonparametric Noise Estimator: We replace our ResNet-based step calibrator with the nonparametric noise estimator [Chen2015-ij] used in RRD [Chung2023-qt]. We determine the step number as the closest cumulative noise to the predicted noise level from the nonparametric noise estimator.
@@
No Augmentation Step Calibrator: We use a step calibrator that is trained without augmentations (i.e. Algorithm C.1 without the red parts).
@@ D.2 Unpaired Images Evaluation additional qualitative results
[width=]figs/supp.unpaired.pdf
D.3 Near-registered dataset collection details
Imaging with low laser power is one of the known ways to produce noisy images. We generate a paired set of near-registered SRH images by imaging a surgical specimen immediately after the SRH imaging system is started (without laser warm-up and optimization), and then image the same specimen after 20 minutes, once the laser has been optimized and sufficiently warmed up. Next, image registration is required to spatially align the low-quality and high-quality images because the tissue specimen may be shifted due to gravity between the two image acquisitions. Paired images were registered by fitting an affine transformation using scale-invariant feature transform [Lowe2004-gq] and RANSAC [derpanis2010overview].
@@ D.3 Near-registered dataset collection details
D.4 Hyperparameters
We list all architecture/training/inference hyperparameters for all components of RSCD in Table 1.
[…]
Base Model | ResNet-50 | |||
@@ D.4 Hyperparameters
Learning Rate | 0.00001 | ||||
Optimizer | Adam | ||||
Epochs | 500 | ||||
Batch Size | 150 | ||||
[…]
Model | UNet | |||
In channels | 2 | ||||
Init dim | 128 | ||||
Out channels | 2 | ||||
Dim mult | [1,2,2,4] | ||||
Resnet block groups | 8 | ||||
Convnext Mult | 2 | ||||
[…]
Noise Schedule | Cosine (s=0.008) | |||
Total Steps (T) | 1000 | ||||
200 | ||||
Batch Size | 8 | ||||
Learning Rate | 0.00001 | ||||
Loss type | L1 | ||||
@@
Optimizer | Adam | ||||
[…]
D.5 Computing Resource and Inference Efficiency
@@
D.6 Human Expert Preference Details
@@
[width=]figs/interface.pdf
@@ E.1 Deep learning-based tumor diagnosis
A portion of our SRH images have ground truth labels on whether the tissue is normal or tumor. Therefore, we took the intersection of that labeled portion and the 12K unpaired low-quality images used for the experiment in section LABEL:sec:unpaired as the dataset for deep learning-based tumor diagnosis. The dataset contains 8,372 low-quality images, each with its ground truth label of tumor/normal.
@@
We used the segmentation model from [Hollon2020-ez], the most influential and widely accepted study on deep learning-based brain tumor diagnosis through SRH images, as our classification model. The model classifies each SRH image into one of three classes: normal, tumor, and non-diagnostic. The model was trained to classify an SRH image as non-diagnostic if the model is unable to extract sufficient information from the image to make a diagnosis, and there are several possibilities for an image to be non-diagnostic, including: (1) the image scanning an irrelevant part of the slide, or (2) the image quality is too low.
@@
As shown in Table LABEL:fig:classifier, the restored images have a significantly higher “correct” rate while having a much lower percentage of misclassified or non-diagnostic images. This means that RSCD can restore a significant portion of the images that were originally classified as non-diagnostic or wrongly classified due to poor quality, and make them high-quality enough to be correctly classified.

### v4

# Step-Calibrated Diffusion for Biomedical Optical Image Restoration

@@ A.1 Image Denoising and Restoration
Image denoising and restoration belong to a set of problems called image inverse problems (Ongie et al. 2020). There are several different levels of difficulty for image denoising and restoration problems. The easiest level is when the corruption process is known as a linear process (such as blurring, inpainting, uniform Gaussian noise, etc) and can be programmatically replicated onto clean images, and the goal is to create an inverse process to the corruption. There have been extensive studies of restoring known corruptions, including supervised approaches (Ongie et al. 2020) and unsupervised approaches (Kawar et al. 2022; Kadkhodaie and Simoncelli 2021; Wang, Yu, and Zhang 2022). The second level of difficulty is when the corruption process is not reproducible, but there are existing pairs of clean and corrupted images allowing for supervised training, and the typical solution is training models to generate the clean image given the corrupted images in a supervised way, such as UNet (Ronneberger, Fischer, and Brox 2015; Manifold et al. 2019) or Conditional Diffusion (Choi et al. 2021; Kawar et al. 2022; Saharia et al. 2022). The most difficult setting is when we don’t have paired data and also cannot reproduce the corruption programmatically; instead, we only have access to unpaired clean images and corrupted images. Existing approaches for this challenging setting include cycle-consistency-based approaches such as CycleGAN (Zhu et al. 2017) and UnitDDPM (Sasaki, Willcocks, and Breckon 2021), VAE-based approaches (Zheng et al. 2023), and flow-based approaches (Du, Chen, and Yang 2020). In addition to the above settings, there are also denoising methods that do not require a training dataset, such as MedianBlur or Deep Image Prior (Ulyanov, Vedaldi, and Lempitsky 2016). General Diffusion Prior (Fei et al. 2023) is a recent unsupervised diffusion-based method that can perform unpaired non-linear blind image restoration, but it is extremely computationally costly as it requires full generative diffusion with multiple gradient steps within each reverse diffusion step.
Restoring noisy SRH images generally falls within the third difficulty level (i.e. no paired data and not a linear inverse problem), as the distribution of noise is unpredictable. Although it is possible to apply supervised methods from the second level by training to remove Gaussian and Laser noises (such as our UNet and Conditional Diffusion baselines), they are generally outperformed by unpaired methods as shown in our experiments. Several recent approaches (including (Chung, Sim, and Ye 2022), (Chung, Lee, and Ye 2023), and (He et al. 2023)) also used the last part of the reverse diffusion process for unpaired image restoration. We discussed these closely related works in more detail in Section Related works.
@@ A.2 Deep Learning applications on SRH
Deep learning techniques have been playing an important role ever since SRH was used for intra-operative neurosurgery, as the very first paper that introduces this optical diagnostic technique (Orringer et al. 2017) trained a multi-layer perceptron to automatically predicts brain tumor subtypes from quantified SRH image attributes. Later, Hollon et al. (Hollon et al. 2020) proposed a CNN-based model that can directly interpret raw SRH images (instead of quantified attributes) and make diagnosis at over 94% accuracy. The first public dataset of clinical SRH images, OpenSRH (Jiang et al. 2022), was released in 2022, and it contains SRH images from 300+ brain tumor patients and 1300+ unique whole slide optical images. DeepGlioma (Hollon et al. 2023) has allowed deep-learning-based models to perform molecular classification of diffuse glioma types via a multimodal model that takes in both SRH images and diffuse glioma genomic data. HiDisc (Jiang et al. 2023) and S3L (Hou et al. 2024) are self-supervised visual representation learning methods that were developed to classify whole-slide SRH images.
All the above works focused on the classification of SRH images, and there have been few previous works that attempted to tackle the unavoidable noise problem of SRH images through image restoration. To the best of our knowledge, the only previous work on this topic was (Manifold et al. 2019), in which the authors collected a paired dataset using low and high laser energy, and trained a U-Net to perform the denoising. In comparison, our method works for all types of noisy SRH images (not just those resulting from low laser power), and we have shown that our method outperforms their U-Net approach.
@@ Appendix C Methodology Additional Details
C.1 Automated Data Filtering
Similar to large language model data curation strategies (Brown et al. 2020), we used a classifier-based filter method to obtain the full high-quality diffusion model training set. A binary classifier was trained on the 4.5K annotated high-quality patches as positive examples (negative examples were randomly selected from the remaining uncurated SRH patch dataset). The trained classifier was then used to split the full SRH dataset into low-quality/high-quality images, resulting in a total of 840K high-quality SRH images for training.
C.2 Step Calibrator Training and Calibration Distribution
We show the full algorithm of training the step calibrator with manually added Gaussian noise and augmentations in Algorithm 2, as well as one example of the augmentation in Figure 11.
@@ C.2 Step Calibrator Training and Calibration Distribution
C.3 Diffusion Model Training
We show the full training algorithm in Algorithm 3 below. It is almost identical to the training algorithm from (Ho, Jain, and Abbeel 2020), except that we only sample up to .
@@ Appendix D Experiment details and additional results
Due to the scale of the experiments, all methods are only run once with one fixed random seed across all images.
@@ D.1 Baseline and Ablation details
CycleGAN (Zhu et al. 2017): We trained a CycleGAN model for image-to-image translation between unpaired low-quality and high-quality images. We used the 4.5K hand-picked high-quality and 3.8K hand-picked low-quality images for training the CyleGAN. Both the high-quality and low-quality image sets contained at least 10 different tumor types. The CycleGAN model performed better when trained on the manually selected images compared to training on the full 840K filtered high-quality images.
Synthetic noise(Lehtinen et al. 2018; Manifold et al. 2019): A U-Net model is trained by adding Gaussian noise to high-quality images to generate synthetic low-quality images. We train only with the 4.5K manually selected high-quality images because we found that the model performs better compared to training on the full 840K filtered high-quality images.
Conditional Diffusion (Kawar et al. 2022; Saharia et al. 2022): We train a conditional generative diffusion model, where we sample from a random Gaussian prior, and generate the restored image conditioned on the low-quality image. The conditioning method is concatenation (i.e. the input image and the conditioning image is concatenated as input to the UNet). We train the model with the 840K HQ training images and fine-tune it with the 4.5K hand-picked images. During training, we add a random amount of Gaussian noise to the high-quality images as conditioning images. We show the full training algorithm in Algorithm.
Regularized Reverse Diffusion (RRD) (Chung, Lee, and Ye 2023) is a recent method developed for restoring noisy medical images. It also uses part of the reverse diffusion process as restoration, but it employs a non-parametric noise estimator (Chen, Zhu, and Heng 2015) to determine the step numbers, and uses a Fourier-feature-based regularization to reduce hallucination. Since we were unable to find any official code for the RRD paper and we deal with images in a vastly different domain from theirs (RRD was originally designed for denoising single-channel MRI images), we trained our own diffusion model following the same process as (Ho, Jain, and Abbeel 2020) and using the same cosine noise schedule as RSCD.
CCDF (Chung, Sim, and Ye 2022) is a recent method for efficient diffusion-based image editing and denoising. CCDF first adds a certain number of steps of Gaussian noise to the input image, and then perform reverse diffusion to create the output image. The number of steps is a fixed hyperparameter, and we follow the original CCDF paper to pick a fixed step number of 20. We trained our own diffusion model following the same process as (Ho, Jain, and Abbeel 2020) and using the same cosine noise schedule as RSCD.
Deep Image Prior (Ulyanov, Vedaldi, and Lempitsky 2016) is an unsupervised image restoration method that works by training a ResNet to reconstruct the noisy image but takes the generation of the model before its convergence as the denoised image. We take the output of step 1800 as our baseline, as images generated from this training step tend to have the best quality.
@@
No Dynamic Recalibration: RSCD without Dynamic Recalibration, i.e. Algorithm 1 without the blue parts. We take the initial step calibration and directly run that many steps of denoising to the end.
Nonparametric Noise Estimator: We replace our ResNet-based step calibrator with the nonparametric noise estimator (Chen, Zhu, and Heng 2015) used in RRD (Chung, Lee, and Ye 2023). We determine the step number as the closest cumulative noise to the predicted noise level from the nonparametric noise estimator.
@@
No Augmentation Step Calibrator: We use a step calibrator that is trained without augmentations (i.e. Algorithm 2 without the red parts).
@@ D.2 Unpaired Images Evaluation additional qualitative results
D.3 Near-registered data collection details
Imaging with low laser power is one of the known ways to produce noisy images. We generate a paired set of near-registered SRH images by imaging a surgical specimen immediately after the SRH imaging system is started (without laser warm-up and optimization), and then image the same specimen after 20 minutes, once the laser has been optimized and sufficiently warmed up. Next, image registration is required to spatially align the low-quality and high-quality images because the tissue specimen may be shifted due to gravity between the two image acquisitions. Paired images were registered by fitting an affine transformation using scale-invariant feature transform (Lowe 2004) and RANSAC (Derpanis 2010).
@@ D.3 Near-registered data collection details
D.4 Near-registered evaluation additional examples
We show additional examples of comparisons between the original/restored/near-registered images from the near-registered experiment in Figure 14.
FID | CMMD | |
(ours) | 32.02 | 0.128 |
32.30 | 0.132 | |
32.64 | 0.126 |
D.5 Hyperparameters
We list all architecture/training/inference hyperparameters for all components of RSCD in Table 4. The number of training epochs are determined based on convergence of loss. Learning rate was chosen between to for optimal training efficiency and performance. Batch size was limited by GPU memory.
Step Calibrator Training | Base Model | ResNet-50 | ||
@@ D.5 Hyperparameters
Learning Rate | 0.00001 | |||
Optimizer | Adam | |||
Epochs | 500 | |||
Batch Size | 150 | |||
Diffusion Model Architecture | Model | UNet | ||
In channels | 2 | |||
Init dim | 128 | |||
Out channels | 2 | |||
Dim mult | [1,2,2,4] | |||
Resnet block groups | 8 | |||
Convnext Mult | 2 | |||
Diffusion Model Training | Noise Schedule | Cosine (s=0.008) | ||
Total Steps (T) | 1000 | |||
200 | |||
Batch Size | 8 | |||
Learning Rate | 0.00001 | |||
Loss type | L1 | |||
@@
Optimizer | Adam | |||
Dynamic Recalibration | Recalibration interval (d) | 10 |
D.6 Computing Resource and Inference Efficiency
@@
D.7 Human Expert Preference Details
@@
D.8 Additional Ablation on Recalibration Interval
We performed additional ablation studies to explore the effect of different recalibration intervals () on the unpaired low-quality images. The results are shown in Table 3. We found that when we change from 10 to 5 or 20, the difference in performance of model is negligible compared to differences between RSCD and other baselines/ablations in Table 1.
@@ E.1 Deep learning-based tumor diagnosis
A portion of our SRH images have ground truth labels on whether the tissue is normal or tumor. Therefore, we took the intersection of that labeled portion and the 12K unpaired low-quality images used for the experiment in section Evaluation on Unpaired Images as the dataset for deep learning-based tumor diagnosis. The dataset contains 8,372 low-quality images, each with its ground truth label of tumor/normal.
@@
We used the segmentation model from (Hollon et al. 2020), the most influential and widely accepted study on deep learning-based brain tumor diagnosis through SRH images, as our classification model. The model classifies each SRH image into one of three classes: normal, tumor, and non-diagnostic. The model was trained to classify an SRH image as non-diagnostic if the model is unable to extract sufficient information from the image to make a diagnosis, and there are several possibilities for an image to be non-diagnostic, including: (1) the image scanning an irrelevant part of the slide, or (2) the image quality is too low.
@@
As shown in Table 7, the restored images have a significantly higher “correct” rate while having a much lower percentage of misclassified or non-diagnostic images. This means that RSCD can restore a significant portion of the images that were originally classified as non-diagnostic or wrongly classified due to poor quality, and make them high-quality enough to be correctly classified.